# Extremal metrics involving scalar curvature

### v3

@@ 4 Proofs of Main Theorems
$\begin{split}&R^{2}_{g_{s}}\cdot g_{s}\\ &=\left(R_{g}+\frac{n-1}{n}s\Delta_{g}R_{g}-s\frac{R_{g}^{2}}{n}+\frac{s}{2}\Delta_{g}R_{g}+s\|\mathrm{Ric}_{g}\|^{2}_{g}+O(s^{2})\right)^{2}\left(g-s\left(\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right)\right)\\ &=R^{2}_{g}\cdot g+s\left(\frac{3n-2}{n}R_{g}\,\Delta_{g}R_{g}+2R_{g}\,\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{R_{g}^{3}}{n}\right)\cdot g-sR^{2}_{g}\cdot\mathrm{Ric}_{g}+O(s^{2})\end{split}$
on $\Lambda^{2}TM$. Hence the desired assertion follows directly from this formula. ∎
Remark 4.1.
From (17), there is a continuous function $F$ on $M$, which is determined by $M,n$ and $g$ such that one can show (2) of Theorem 1.3 on any compact subset $K\subset M$ provided that
@@
$\inf\left\{s>0:~\|\overset{\circ}{\mathrm{Ric}}_{g}\|_{g}^{2}-R_{g}\left(\overline{\lambda}_{Ric}-\frac{R_{g}}{n}\right)>sF~\mathrm{on}~K\right\}>0.$ (18)
In particular, if the minimum of
$\|\overset{\circ}{\mathrm{Ric}}_{g}\|_{g}^{2}-R_{g}\left(\overline{\lambda}_{Ric}-\frac{R_{g}}{n}\right)$ (19)
on $K$ is positive, then the above condition (18) holds. For any compact set $K\subset\{x\in M\mid~\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}(x)\neq 0\}$, the condition: $R_{g}\equiv\mathrm{const}\leq 0$ on $K$, especially implies (19). Similarly, there is a continuous symmetric $(0,2)$-tensor $h$ on $M$, which is determined by $M,n$ and $g$ such that one can show (7) of Theorem 1.5 on any compact subset $K\subset M$ provided that
@@
$\inf\left\{s>0:~\left(\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}^{2}+\frac{R^{2}_{g}}{n}\right)g-R_{g}\cdot\mathrm{Ric}_{g}>sh~\mathrm{on}~K\right\}>0.$ (20)
In particular, the condition: $R_{g}\equiv\mathrm{const}\leq 0$ on $K$ and $\min_{K}\underline{\lambda}_{Ric}\geq\frac{R_{g}}{n}$, especially implies (20).
@@ 5.1 On the product of two Einstein spaces
Let $n,m\geq 2$ and $(\mathbb{S}^{n},g_{\mathbb{S}^{n}})$, $(\mathbb{H}^{m},g_{\mathbb{H}^{m}})$ be space forms with sectional curvature $1$ and 1-1 respectively. Consider the metric of the form
$g_{\lambda}:=g_{\mathbb{S}^{n}}+\lambda g_{\mathbb{H}^{m}},$
where $\lambda\geq 1$ is a scaling constant of the hyperbolic space. Then, for each point $x\in\mathbb{S}^{n}\times\mathbb{H}^{m}$,
$\mathrm{Ric}(g_{\lambda})=\begin{cases}(n-1)g_{\mathbb{S}^{n}}&\mathrm{on}~T_{\mathbf{p}_{1}x}\mathbb{S}^{n}\subset T_{x}(\mathbb{S}^{n}\times\mathbb{H}^{m}),\\ -(m-1)g_{\mathbb{H}^{m}}&\mathrm{on}~T_{\mathbf{p}_{2}x}\mathbb{H}^{m}\subset T_{x}(\mathbb{S}^{n}\times\mathbb{H}^{m}),\end{cases}$
@@
$R(g_{\lambda})=n(n-1)-\frac{m(m-1)}{\lambda}.$
Here, $\mathbf{p}_{i}~(i=1,2)$ denotes the natural projection from respectively $\mathbb{S}^{n}$ and $\mathbb{H}^{m}$ to $\mathbb{S}^{n}\times\mathbb{H}^{m}$.
So, if we put $n=m$, then one can check that $R(g_{\lambda})\geq 0$ and $g_{\lambda}$ satisfies the assumptions of Theorems 1.3, 1.4, 1.5 and 1.6 for all $\lambda\geq 1$. Moreover, if we put $n=m+1$, one can also check that $R(g_{1})\geq 0$ and $g_{1}$ satisfies the assumptions of Theorems 1.3, 1.4, 1.5 and 1.6. On the other hand, we can directly check that $g_{1}$ is none of type I, II, III or IV scalar curvature rigid in the sense of Listing. In fact, the family $\{g_{\lambda}\}_{\lambda\geq 1}$ gives a deformation which suggests that $g_{1}$ is not rigid in each sense. In particular, for every dimension $n\geq 4$, there is an $n$-dimensional manifold on which there is a non-rigid metric $g$ in each sense.
@@ 5.2 Examples on Lie groups
We can construct examples of left invariant metrics on some three dimensional Lie groups that satisfy the assumptions of our main theorems in Section 1. Curvatures of left invariant metrics on Lie groups have been studied by Milnor in [27]. Especially, the case of three-dimensional unimodular is written in Section 4 of [27]. The following are examples of metrics on unimodular three-dimensional Lie groups, which satisfy the assumption of Theorem 1.3 or 1.5.
@@
    Hence, if $0<\varepsilon<4$, then $g_{s,t}$ has a positive constant scalar curvature and satisfies the assumptions of Theorems 1.3, 1.4, 1.5 and 1.6. Moreover, from [14, Example 2], if $\varepsilon\leq 2$ (resp. $\varepsilon<2$), then $g_{s,t}$ is a (resp. unique) Yamabe metric with positive scalar curvature. On the other hand, if $\varepsilon\leq 0$, then $g_{s,t}$ has a non-positive constant scalar curvature and satisfies the assumptions of Theorems 1.3, 1.4 and 1.5 but does not one of Theorem 1.6.
@@ 5.3 Examples on the total spaces of Riemannian submersions with totally geodesic fibers
  • ([26, Section 5]) Let $\pi:S^{4n+3}\rightarrow\mathbb{H}P^{n}$ be the Hopf fibration whose fibers are the standard unit $3$-sphere $\mathbb{S}^{3}$. We denote the scalar curvatures of the fibers, the base space and the total space by $R^{F},R^{B}$ and $R^{M}$, respectively. Then,
@@
    $\begin{split}\mathrm{Ric}^{t}(U,V)&=\left(\frac{R^{F}}{t\,\mathrm{dim}F}+t\left(\frac{R^{M}}{\mathrm{dim}F+\mathrm{dim}B}-\frac{R^{F}}{\mathrm{dim}F}\right)\right)\,g_{t}(U,V),\\ \mathrm{Ric}^{t}(X,Y)&=\left(\frac{R^{B}}{t\,\mathrm{dim}B}+t\left(\frac{R^{M}}{\mathrm{dim}F+\mathrm{dim}B}-\frac{R^{B}}{\mathrm{dim}B}\right)\right)\,g_{t}(X,Y),\end{split}$
@@
    $\begin{split}\|\overset{\circ}{\mathrm{Ric}_{g_{t}}}\|^{2}_{g_{t}}&=3\left(\frac{2}{t}+4nt-\frac{1}{4n+3}\left(\frac{6}{t}-12nt+16n(n+2)\right)\right)^{2}\\ &+4n\left(4(n+2)-6t-\frac{1}{4n+3}\left(\frac{6}{t}-12nt+16n(n+2)\right)\right)^{2}\\ &=3\left(\frac{8n}{4n+3}\cdot\frac{1}{t}+\frac{4n(4n+6)}{4n+3}t-\frac{16n(n+2)}{4n+3}\right)^{2}\\ &+4n\left(\frac{6}{4n+3}\cdot\frac{1}{t}+\frac{12n+18}{4n+3}t-\frac{4(n+2)(4n-1)}{4n+3}\right)^{2}.\end{split}$
@@
    $\begin{cases}\lambda_{Ric_{t}}^{V}\geq\lambda_{Ric_{t}}^{V}&0<t\leq\frac{1}{2n+3}~\mathrm{or}~t\geq 1\\ \lambda_{Ric_{t}}^{H}\geq\lambda_{Ric_{t}}^{H}&\frac{1}{2n+3}\leq t\leq 1.\end{cases}$
@@
    $\begin{split}&R_{g_{t}}\left(\lambda_{Ric_{t}}^{V}-\frac{R_{g_{t}}}{\mathrm{dim}M}\right)\\ &=\left(\frac{6}{t}-12nt+16n(n+2)\right)\left(\frac{8n}{4n+3}\cdot\frac{1}{t}+\frac{4n(4n+6)}{4n+3}t-\frac{16n(n+2)}{4n+3}\right),\\ &R_{g_{t}}\left(\lambda_{Ric_{t}}^{H}-\frac{R_{g_{t}}}{\mathrm{dim}M}\right)\\ &=\left(\frac{6}{t}-12nt+16n(n+2)\right)\left(\frac{6}{4n+3}\cdot\frac{1}{t}+\frac{12n+18}{4n+3}t-\frac{4(n+2)(4n-1)}{4n+3}\right).\end{split}$
    Hence, if $0<t<<1$ or $t>>1$, then
    $\|\overset{\circ}{\mathrm{Ric}_{g_{t}}}\|_{g_{t}}^{2}\cdot g_{t}>R_{g_{t}}\left(\mathrm{Ric}_{g_{t}}-\frac{R_{g_{t}}}{\mathrm{dim}M}\cdot g_{t}\right).$
    Therefore, $(\mathbb{S}^{4n+3},g_{t})$ satisfies the assumption of Theorems 1.3, 1.4 and 1.5 for all $0<t<<1$ and $t>>1$. Moreover, since $R_{g_{t}}>0$ on $M$, $g_{t}$ also satisfies the assumption of Theorem 1.6 for all $0<t<<1$ and $t>>1$.
  • ([26, Section 5]) Consider the fibration $\mathbb{C}P^{2n+1}\rightarrow\mathbb{H}P^{n}$ whose fibers are $S^{2}(4)$ with constant sectional curvature $4$. Then
@@
    $\begin{cases}\lambda_{Ric_{t}}^{V}\geq\lambda_{Ric_{t}}^{V}&0<t\leq\frac{1}{n+1}~\mathrm{or}~t\geq 1\\ \lambda_{Ric_{t}}^{H}\geq\lambda_{Ric_{t}}^{H}&\frac{1}{n+1}\leq t\leq 1.\end{cases}$
    Matsuzawa [26] observed that the canonical variation $g_{t}$ is an Einstein metric on $\mathbb{S}^{4n+3}$ if and only if $t=1$ or $t=\frac{1}{n+1}$. Then, we have
@@
    Hence, if $0<t<<1$ or $t>>1$, then
    $\|\overset{\circ}{\mathrm{Ric}_{g_{t}}}\|_{g_{t}}^{2}\cdot g_{t}>R_{g_{t}}\left(\mathrm{Ric}_{g_{t}}-\frac{R_{g_{t}}}{\mathrm{dim}M}\cdot g_{t}\right).$
Therefore, $(\mathbb{C}P^{2n+1},g_{t})$ satisfies the assumption of Theorems 1.3, 1.4 and 1.5 for all $0<t<<1$ and $t>>1$. Moreover, since $R_{g_{t}}>0$ on $M$, $g_{t}$ also satisfies the assumption of Theorem 1.6 for all $0<t<<1$ and $t>>1$.

### v4

@@ 4 Proofs of Main Theorems
$\begin{split}R_{g_{s}}&-R_{g}\cdot\|g\|^{2}_{1,\,g_{s}}\\ &=R_{g}+se^{f}\left[\left\langle\mathrm{Hess}f+\nabla f\otimes\nabla f,\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right\rangle_{g}-\|\mathrm{Ric}_{g}\|_{g}^{2}\right.\\ &\qquad\left.+\frac{R^{2}_{g}}{n}+\left(1-\frac{1}{n}\right)\langle\nabla f,\nabla R_{g}\rangle_{g}+\left(\frac{1}{2}-\frac{1}{n}\right)\Delta_{g}R_{g}\right]+O(s^{2})\\ &\qquad-R_{g}\,\frac{1}{1+s\min_{v\in T_{x}M,\,|v|_{g}=1}e^{f}\overset{\circ}{\mathrm{Ric}_{g}}(v,v)}\\ &=R_{g}+se^{f}\left[\left\langle\mathrm{Hess}f+\nabla f\otimes\nabla f,\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right\rangle_{g}-\|\mathrm{Ric}_{g}\|_{g}^{2}\right.\\ &\qquad\left.+\frac{R^{2}_{g}}{n}+\left(1-\frac{1}{n}\right)\langle\nabla f,\nabla R_{g}\rangle_{g}+\left(\frac{1}{2}-\frac{1}{n}\right)\Delta_{g}R_{g}\right]+O(s^{2})\\ &\qquad-R_{g}\left(1-s\min_{v\in T_{x}M,\,|v|_{g}=1}e^{f}\overset{\circ}{\mathrm{Ric}_{g}}(v,v)+O(s^{2})\right).\end{split}$
Then the assertion immediately follows from this. ∎
Proof of Theorem 1.6.
Let $g_{s}=g+s\,e^{f}\,\overset{\circ}{\mathrm{Ric}_{g}}$ Then, as in the above proof of Theorem 1.5, we have
$\begin{split}R_{g_{s}}&-R_{g}\cdot\|g\|_{2,g_{s}}\\ &=R_{g}+se^{f}\left[\left\langle\mathrm{Hess}f+\nabla f\otimes\nabla f,\mathrm{Ric}_{g}-\frac{R_{g}}{n}g\right\rangle_{g}-\|\mathrm{Ric}_{g}\|_{g}^{2}\right.\\ &\qquad\left.+\frac{R^{2}_{g}}{n}+\left(1-\frac{1}{n}\right)\langle\nabla f,\nabla R_{g}\rangle_{g}+\left(\frac{1}{2}-\frac{1}{n}\right)\Delta_{g}R_{g}\right]+O(s^{2})\\ &\qquad-R_{g}\,\sqrt{\max_{v\in\Lambda^{2}T_{x}M\setminus\{0\}}\frac{g(v,v)}{\left(g+s\,e^{f}\,\overset{\circ}{\mathrm{Ric}_{g}}\right)(v,v)}}.\end{split}$ (18)
Here, for any regular $n\times n$ matrix AA and a $n\times n$ matrix $B$,
$\det{(A+tB)}=\det{A}\cdot\det{\left(\mathrm{Id}_{n}+tA^{-1}B\right)}=\det{A}\left(1+t\,\mathrm{trace}(A^{-1}B)+O(t^{2})\right).$
So, we have
@@
$\begin{split}\sqrt{\max_{v\in\Lambda^{2}T_{x}M\setminus\{0\}}\frac{g(v,v)}{\left(g+s\,e^{f}\,\overset{\circ}{\mathrm{Ric}_{g}}\right)(v,v)}}&=1-\frac{s}{2}\min_{v\in\Lambda^{2}T_{x}M\setminus\{0\}}e^{f}g^{-1}\overset{\circ}{\mathrm{Ric}_{g}}(v,v)+O(s^{2}).\end{split}$ (19)
Putting (18) and (19) together, we can immediately get the assertion. ∎
Proof of Theorem 1.8.
As the proof of Theorem 1.6, one can also compute the difference $R_{g_{s}}^{2}\cdot g-R_{g}^{2}\cdot g$ on $\Lambda^{2}TM$. Then, from the above calculation in the proof of Theorem 1.6, it is easy to see that the assertion holds . ∎
As we can see from some of the following examples, there is an example of Yamabe metric (hence it is semi-stable csc) but not type III scalar curvature rigid in the sense of Listing (Definition 1.3). More abstractly, we can also observe the following. According to [1], if $g$ is a strongly stable unique non-negative Yamabe metric with unit volume in its conformal class (then every metric sufficiently $C^{\infty}$-close to $g$ also contains the unique Yamabe metric in its conformal class), then for any $(0,2)$-tensor $h$ with $\mathrm{tr}_{g}h=0$, we have
for all sufficiently small $t>0$. Here, $\gamma_{t}\in[g_{t}:=g+th]$ is the unique Yamabe metric in its conformal class with unit volume. Since $\gamma_{t}$ is the unique Yamabe (hence constant scalar curvature) metric in $\gamma_{t}]$ for all sufficiently small $t>0$,
$\gamma_{t}-g=t\,\mathrm{pr}_{TT}(h)+o(t)=th+o(t),$
where $\mathrm{pr}_{TT}$ is the projection onto the subspace consisting of all tt-tensors in $T_{g}\mathcal{M}_{1}$. Here, $\mathcal{M}_{1}$ is the space of all Riemannain metric on $M$ with unit volume and $T_{g}\mathcal{M}_{1}$ is the tangent space of $\mathcal{M}_{1}$ at the metric $g$. Therefore, if we take $h=-\overset{\circ}{\mathrm{Ric}}_{g}$, then we have
@@
$\begin{split}R_{\gamma_{t}}\cdot\gamma_{t}-R_{g}\cdot g&=t\left(\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{L^{2}(M,g)}g-R_{g}\cdot\overset{\circ}{\mathrm{Ric}}_{g}\right)+o(t)\\ &=t\left(\left(\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{L^{2}(M,g)}+\frac{R^{2}_{g}}{n}\right)g-R_{g}\cdot\mathrm{Ric}_{g}\right)+o(t).\end{split}$
By [25, Corollary 2], for every closed manifold $M$ of dimension $\geq 3$ with positive Yamabe invariant, there is a non-Ricci-flat scalar-flat metric $g$ on $M$. Thus, the above argument with $h=\mathrm{Ric}_{g}$ for such $g$ implies that $R_{\gamma_{t}}>0$ for any sufficiently small $t>0$ where $\gamma_{t}\in[g+t\,\mathrm{Ric}_{g}]$ is the unique unit-volume Yamabe metric in its conformal class. Therefore, such a metric $g$ is not a local maximizer of the functional : $\mathcal{M}_{1}\ni h\mapsto Y(M,[h])$ and not type III scalar curvature rigid in the sense of Listing.
Question 4.1.
@@ 5.1 On the product of two Einstein spaces
Let $m,n\geq 2$ and $(M^{m},g_{M})$, $(N^{n},g_{N})$ be closed Riemannian manifolds with constant curvature $1$ and 1-1 respectively. Consider the metric of the form
$g_{\lambda}:=g_{M}+\lambda g_{N},$
where $\lambda\geq 1$ is a scaling constant. Then, for each point $x\in M\times N$,
$\mathrm{Ric}(g_{\lambda})=\begin{cases}(m-1)g_{M^{m}}&\mathrm{on}~T_{\mathbf{p}_{1}x}M^{m}\subset T_{x}(M^{m}\times N^{n}),\\ -(n-1)g_{N^{n}}&\mathrm{on}~T_{\mathbf{p}_{2}x}N^{n}\subset T_{x}(M^{m}\times N^{n}),\end{cases}$
@@
$R(g_{\lambda})=m(m-1)-\frac{n(n-1)}{\lambda}.$
Here, $\mathbf{p}_{i}~(i=1,2)$ denotes the natural projection from, respectively $M^{m}$ and $N^{n}$ to $M^{n}\times N^{n}$.
So, if we put $n=m$, then one can check that $R(g_{\lambda})\geq 0$ and $g_{\lambda}$ satisfies the assumptions of Theorems 1.5, 1.6, 1.7 and 1.8 for all $\lambda\geq 1$. Moreover, if we put $m=n+1$, one can also check that $R(g_{1})\geq 0$ and $g_{1}$ satisfies the assumptions of Theorems 1.5, 1.6, 1.7 and 1.8. On the other hand, we can directly check that $g_{1}$ is none of type I, II, III or IV scalar curvature rigid in the sense of Listing. In fact, the family $\{g_{\lambda}\}_{\lambda\geq 1}$ gives a deformation which suggests that $g_{1}$ is not rigid in each sense. In particular, for every dimension $n\geq 4$, there is an $n$-dimensional manifold on which there is a non-rigid metric $g$ in each sense.
@@ 5.2 Examples on Lie groups
We can construct examples of left invariant metrics on some three dimensional Lie groups that satisfy the assumptions of our main theorems in Section 1. Curvatures of left invariant metrics on Lie groups have been studied by Milnor in [27]. Especially, the case of three-dimensional unimodular is written in Section 4 of [27]. The following are examples of metrics on unimodular three-dimensional Lie groups, which satisfy the assumption of Theorem 1.5 or 1.7.
@@
    Hence, if $0<\varepsilon<4$, then $g_{s,t}$ has a positive constant scalar curvature and satisfies the assumptions of Theorems 1.5, 1.6, 1.7 and 1.8. Moreover, from [14, Example 2], if $\varepsilon\leq 2$ (resp. $\varepsilon<2$), then $g_{s,t}$ is a (resp. unique) Yamabe metric with positive scalar curvature. On the other hand, if $\varepsilon\leq 0$, then $g_{s,t}$ has a non-positive constant scalar curvature and satisfies the assumptions of Theorems 1.5, 1.6 and 1.7 but does not one of Theorem 1.8.
@@ 5.3 Examples on the total spaces of Riemannian submersions with totally geodesic fibers
  • ([26, Section 5] or [4, 9.82 Example 2]) Let $\pi:S^{4n+3}\rightarrow\mathbb{H}P^{n}$ be the Hopf fibration whose fibers are the standard unit $3$-sphere $\mathbb{S}^{3}$. We denote the scalar curvatures of the fibers, the base space and the total space by $R^{F},R^{B}$ and $R^{M}$, respectively. Then,
@@
    $\begin{split}\mathrm{Ric}^{t}(U,V)&=\left(\frac{R^{F}}{t\,\mathrm{dim}F}+t\left(\frac{R^{M}}{\mathrm{dim}F+\mathrm{dim}B}-\frac{R^{F}}{\mathrm{dim}F}\right)\right)\,g_{t}(U,V),\\ \mathrm{Ric}^{t}(X,Y)&=\left(\frac{R^{B}}{\mathrm{dim}B}+t\left(\frac{R^{M}}{\mathrm{dim}F+\mathrm{dim}B}-\frac{R^{B}}{\mathrm{dim}B}\right)\right)\,g_{t}(X,Y),\end{split}$
@@
    $\begin{split}\|\overset{\circ}{\mathrm{Ric}_{g_{t}}}\|^{2}_{g_{t}}&=3\left(\frac{2}{t}+4nt-\frac{1}{4n+3}\left(\frac{6}{t}-12nt+16n(n+2)\right)\right)^{2}\\ &+4n\left(4(n+2)-6t-\frac{1}{4n+3}\left(\frac{6}{t}-12nt+16n(n+2)\right)\right)^{2}\\ &=3\left(\frac{8n}{4n+3}\cdot\frac{1}{t}+\frac{4n(4n+6)}{4n+3}t-\frac{16n(n+2)}{4n+3}\right)^{2}\\ &+4n\left(\frac{6}{4n+3}\cdot\frac{1}{t}+\frac{12n+18}{4n+3}t-\frac{12(n+2)}{4n+3}\right)^{2}.\end{split}$
@@
    $\begin{cases}\lambda_{Ric_{t}}^{V}\geq\lambda_{Ric_{t}}^{H}&0<t\leq\frac{1}{2n+3}~\mathrm{or}~t\geq 1\\ \lambda_{Ric_{t}}^{H}\geq\lambda_{Ric_{t}}^{V}&\frac{1}{2n+3}\leq t\leq 1.\end{cases}$
@@
    $\begin{split}&R_{g_{t}}\left(\lambda_{Ric_{t}}^{V}-\frac{R_{g_{t}}}{\mathrm{dim}M}\right)\\ &=\left(\frac{6}{t}-12nt+16n(n+2)\right)\left(\frac{8n}{4n+3}\cdot\frac{1}{t}+\frac{4n(4n+6)}{4n+3}t-\frac{16n(n+2)}{4n+3}\right),\\ &R_{g_{t}}\left(\lambda_{Ric_{t}}^{H}-\frac{R_{g_{t}}}{\mathrm{dim}M}\right)\\ &=\left(\frac{6}{t}-12nt+16n(n+2)\right)\left(\frac{6}{4n+3}\cdot\frac{1}{t}+\frac{12n+18}{4n+3}t-\frac{12(n+2)}{4n+3}\right).\end{split}$
    And,
    $\begin{split}&\|\overset{\circ}{\mathrm{Ric}_{g_{t}}}\|^{2}_{g_{t}}-R_{g_{t}}\left(\lambda_{Ric_{t}}^{V}-\frac{R_{g_{t}}}{\mathrm{dim}M}\right)=\frac{16n}{t}((2n+3)t-1)(t-1)(3t-2(n+2)),\\ &\|\overset{\circ}{\mathrm{Ric}_{g_{t}}}\|^{2}_{g_{t}}-R_{g_{t}}\left(\lambda_{Ric_{t}}^{H}-\frac{R_{g_{t}}}{\mathrm{dim}M}\right)\\ &=\frac{12}{(4n+3)t^{2}}((2n+3)t-1)(t-1)\left((8n^{2}+18n)t^{2}-16n(n+2)t+4n-3\right).\end{split}$
    Hence, if $t_{-}<t<\frac{1}{2n+3}$, $\frac{1}{2n+3}<t<1$, $1<t<t_{+}$ or $t>\frac{2}{3}(n+2)$, then
    $\|\overset{\circ}{\mathrm{Ric}_{g_{t}}}\|_{g_{t}}^{2}\cdot g_{t}-R_{g_{t}}\left(\mathrm{Ric}_{g_{t}}-\frac{R_{g_{t}}}{\mathrm{dim}M}\cdot g_{t}\right)$
    is positive or negative definite on $M$. Here, $t_{\pm}$ are the real roots of the equation: $(8n^{2}+18n)t^{2}-16n(n+2)t+4n-3=0$. Note that $0<t_{-}<\frac{1}{2n+3}$ and $1<t_{+}<\frac{2}{3}(n+2)$. Therefore, $(\mathbb{S}^{4n+3},g_{t})$ satisfies the assumption of Theorems 1.5, 1.6, 1.7 and 1.8 for such $(\mathbb{S}^{4n+3},g_{t})$. Moreover, $g_{t}$ is a Yamabe metric if $t\geq\frac{4n+5}{3}$.
  • ([26, Section 5] or [4, 9.83 Example 3]) Consider the fibration $\mathbb{C}P^{2n+1}\rightarrow\mathbb{H}P^{n}$ whose fibers are $S^{2}(4)$ with constant sectional curvature $4$. Then
@@
    $\begin{cases}\lambda_{Ric_{t}}^{V}\geq\lambda_{Ric_{t}}^{H}&0<t\leq\frac{1}{n+1}~\mathrm{or}~t\geq 1\\ \lambda_{Ric_{t}}^{H}\geq\lambda_{Ric_{t}}^{V}&\frac{1}{n+1}\leq t\leq 1.\end{cases}$
    Matsuzawa [26] observed that the canonical variation $g_{t}$ is an Einstein metric on $\mathbb{C}P^{2n+1}$ if and only if $t=1$ or $t=\frac{1}{n+1}$. Then, we have
@@
    And,
    $\begin{split}&\|\overset{\circ}{\mathrm{Ric}_{g_{t}}}\|^{2}_{g_{t}}-R_{g_{t}}\left(\lambda_{Ric_{t}}^{V}-\frac{R_{g_{t}}}{\mathrm{dim}M}\right)=\frac{64n}{t^{2}}(t-n-2)((n+1)t-1)(t-1),\\ &\|\overset{\circ}{\mathrm{Ric}_{g_{t}}}\|^{2}_{g_{t}}-R_{g_{t}}\left(\lambda_{Ric_{t}}^{H}-\frac{R_{g_{t}}}{\mathrm{dim}M}\right)\\ &\quad=\frac{32}{(2n+1)t^{2}}((n+1)t-1)(t-1)\left(n(2n+3)t^{2}-4n(n+2)t+2n-1\right).\end{split}$
    Hence, if $t_{-}<t<\frac{1}{n+1}$, $\frac{1}{n+1}<t<1$, $1<t<t_{+}$, or $t>n+2$, then
    $\|\overset{\circ}{\mathrm{Ric}_{g_{t}}}\|_{g_{t}}^{2}\cdot g_{t}-R_{g_{t}}\left(\mathrm{Ric}_{g_{t}}-\frac{R_{g_{t}}}{\mathrm{dim}M}\cdot g_{t}\right)$
    is positive or negative definite on $M$. Here, $t_{\pm}$ are the real roots of the equation: $n(2n+3)t^{2}-4n(n+2)t+2n-1=0$. Note that $0<t_{-}<\frac{1}{n+1}$ and $1<t_{+}<n+2$. Therefore, $(\mathbb{C}P^{2n+1},g_{t})$ satisfies the assumption of Theorems 1.5, 1.6, 1.7 and 1.8 for $(\mathbb{C}P^{2n+1},g_{t})$. Moreover, $g_{t}$ is a Yamabe metric if $t\geq 2n+3$.
  • ([4, 9.84 Example 4]) Let $\pi:\mathbb{S}^{15}\rightarrow\mathbb{S}^{8}$ be the Hopf fibratioin with fiber $\mathbb{S}^{7}$. Then, the metric $g$ on the total space $\mathbb{S}^{15}$ and $g^{F}$ on the fiber $\mathbb{S}^{7}$ have constant curvature $1$, and the metric $g^{B}$ on the base has constant curvature $4$. Hence, we have
    $R^{F}=R_{g_{\mathbb{S}^{7}(1)}}=42,R^{M}=R_{g_{\mathbb{S}^{15}(1)}}=210,R^{B}=R_{g_{\mathbb{S}^{8}(1/2)}}=4\cdot 56=224.$
    And, let $g_{t}~(t>0)$ be the canonical variation of $g$, then for vertical vectors $U,V$ and horizontal vectors $X,Y$,
    $\begin{split}\mathrm{Ric}_{g_{t}}(U,V)&=\left(\frac{6}{t}+8t\right)\langle U,V\rangle_{t},\\ \mathrm{Ric}_{g_{t}}(X,Y)&=(28-14t)\langle X,Y\rangle_{t}.\end{split}$
    Then, one can calculate as
    $\begin{split}R_{g_{t}}&=\frac{42}{t}-56t+224,\\ \|\overset{\circ}{\mathrm{Ric}_{g_{t}}}\|^{2}_{g_{t}}&=\left(7\left(\frac{16}{15}\right)^{2}+8\left(\frac{14}{15}\right)^{2}\right)\frac{1}{t^{2}}(11t-3)^{2}(t-1)^{2},\\ R_{g_{t}}\left(\lambda^{V}_{t}-\frac{R_{g_{t}}}{15}\right)&=\frac{14\cdot 16}{15t^{2}}(3+6t-4t^{2})(11t-3)(t-1),\\ R_{g_{t}}\left(\lambda^{H}_{t}-\frac{R_{g_{t}}}{15}\right)&=-\frac{14^{2}}{15t^{2}}(3+6t-4t^{2})(11t-3)(t-1).\end{split}$
    Moreover,
    $\begin{cases}\lambda^{V}_{t}\geq\lambda^{H}_{t}&t\leq\frac{3}{11}~\mathrm{or}~t\geq 1\\ \lambda^{H}_{t}\geq\lambda^{V}_{t}&\frac{3}{11}\leq t\leq 1,\end{cases}$
    and $(\mathbb{S}^{15},g_{t})$ is Einstein if and only if $t=\frac{3}{11}$ or $1$. $(\mathbb{S}^{15},g_{t})$ satisfies the assumption of Theorem 1.8 if $0<t<<\frac{3}{11}$ or $t>>1$. Moreover, $g_{t}$ is a Yamabe metric if $t\geq 3$.
All of the manifolds in the above examples have positive Yamabe invariant.